A stock has the following stock information:

Year

Stock Price

2014

$         22.75

2015

$         22.00

2016

$         23.75

2017

$         26.00

What is the arithmetic average return of this stock? ENTER YOUR ANSWER AS A PERCENTAGE WITH ONE DECIMAL PLACE (e.g., 12.1) AND NOT AS A DECIMAL (e.g., 0.121). ROUND TO THE NEAREST TENTH OF A PERCENT. DO NOT USE THE PERCENT SIGN (%) IN YOUR ANSWER.
Answer #1

Average Return in 2015 = (22 - 22.75)/22.75 = -3.30%

Average Return in 2016 = (23.75 - 22)/22 = 7.95%

Average Return in 2017 = (26 - 23.75)/23.75 = 9.47%

Average Growth Rate = (-0.033 + 0.0795 + 0.0947)/3

Average Growth Rate = 4.71%